# Convexity on manifolds without focal points and spectrum of the Laplacian

### v1

@@ 3. A strictly convex function on a manifold without focal points
$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert~{}\mathrm{d}\eta(v)\quad(D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)~{}\text{is a bounded linear map for}~{}v\in S_{p}M).$
@@
The restriction of the continuous map $w\mapsto\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert$ on $SM$ to $S_{p}M$ shows that there exists some $L>0$ such that $\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $w\in S_{p}M$. In particular, $\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $v\in S_{o}M$. So another application of the dominated convergence theorem shows that
${\mathrm{H}}^{F}(p)(w,w)=\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v).$
@@
The convexity of $F$ follows from the convexity of the Busemann functions. We now claim that $F$ is, in fact, strictly convex. Suppose there exists $p\in M$ and $0\neq w\in T_{p}M$ such that ${\mathrm{H}}^{F}(p)(w,w)=0$ then
$\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(w,w)~{}\mathrm{d}\eta(v)=0$
@@
Since $p\mapsto{\mathrm{H}}^{b_{v}}(p)(w,w)$ is a nonnegative continuous function in $v$, we have ${\mathrm{H}}^{b_{v}}(p)(w,w)=0$ for all $v\in S_{o}M$. Using (2.2),
@@
(3.1) ${\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}},{D_{\nabla b_{v}(p)}(0)w^{\perp}\rangle}=0,\quad\text{for all}~{}v\in S_{o}M.$
@@
It is known from [13, Section 5] that for $v\in SM$ and all parallel vector fields $x$ normal to $\gamma_{v}$, $\frac{d}{dt}(\left\lVert D_{v}(t)x\right\rVert^{2})\leq 0$ i.e., the function $f_{v,x}(t)=\frac{1}{2}\left\lVert D_{v}(t)x\right\rVert^{2}$ is $C^{2}$ and monotonically decreasing. From (3.1) we see that for each $v\in S_{o}M$, $f_{\nabla b_{v}(p),w^{\perp}}$ has a critical point at 00. We claim that $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. If not, then 00 is either a local minimum or local maximum. Suppose 00 is a local minimum then there exists a $\delta$-neighbourhood of 00 such that $f_{\nabla b_{v}(p),w^{\perp}}(0)\leq f_{\nabla b_{v}(p),w^{\perp}}(t)$ for all $t\in(-\delta,\delta)$. Since $f$ is monotonically decreasing we have $f_{\nabla b_{v}(p),w^{\perp}}(t)=f_{\nabla b_{v}(p),w^{\perp}}(0)$ for all $t\in(0,\delta)$. Hence $f_{\nabla b_{v}(p),w^{\perp}}$ is constant on $(0,\delta)$. Observe that $f_{\nabla b_{v}(p),w^{\perp}}$ is $C^{2}$ and so $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=\lim\limits_{t\to 0_{+}}f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(t)=0$, a contradiction. A similar argument could be carried out to show that 00 cannot be a local maximum of $f_{\nabla b_{v}(p),w^{\perp}}$. Hence 00 is a degenerate critical point of $f_{\nabla b_{v}(p),w^{\perp}}$ i.e, $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. Then the Jacobi equation for $D_{\nabla b_{v}(p)}$,
@@
00 0 $=f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)$
@@
(3.2) $\implies K(\nabla b_{v}(p),w^{\perp})$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}\geq 0,\quad\text{for all $v\in S_{o}M$}$
@@
Now for $\{x_{1},\cdots,x_{n-1}\}$ orthonormal vectors in $T_{p}M$ such that $x_{i}\perp w$, $Ricc(w,w)=\sum\limits_{i=1}^{n}K(x_{i},w)\geq 0$. This is a contradiction to the hypothesis that $M$ has negative Ricci curvature. Hence ${\mathrm{H}}^{F}(p)(w,w)>0$ for every $p\in M$ and every $0\neq w\in T_{p}M$ proving the strict convexity of $F$. ∎
While there are several geometrically motivated convex functions on a Riemannian manifold, typically they have low regularity. Even in situations where such convex functions have higher regularity they often do not have other useful properties such as having bounded gradient. On the one hand, distance functions and Busemann functions are examples of convex functions with bounded gradient having low regularity. On the other hand, on complete simply connected non-compact Riemannian manifolds without focal points, the squared distance function is a smooth strictly convex function on $M$ [11, Lemma 1] whose gradient is not bounded. It is in this context that we have the following approximation result for strictly convex functions with bounded gradient. Note that the smoothing approximation fails if we drop the assumption of strict convexity [26]. Recall that a $C^{2}$-function is strictly convex if its Hessian is positive definite. It was proved in [17, Theorem 2(a)] that a Lipschitz $C^{2}$-function which is strictly convex admits a smooth strictly convex approximation which is also Lipschitz. From this and using the fact that Lipschitz functions have bounded gradient we have the following corollary to our Theorem 3.1.
Corollary 3.2.
Let $M$ be as in Theorem 3.1. Then for any $C>1$, $M$ admits a smooth strictly convex function $f$ such that $\left\lVert\nabla f\right\rVert\leq C$.
The property of having bounded gradient is useful in applications as can be seen below in Corollary 3.3 and Theorem 4.7.
Let $(M,g,J)$ be a Kähler manifold as in the introduction. The complex analogue of convex functions is given by the notion of plurisubharmonic functions. These are defined using the complex Hessian called the Levi form. In local coordinates, the Levi form of a function $f$ is given by
$Lf=4\sum\limits_{i,j=1}^{n}\frac{\partial^{2}f}{\partial z_{i}\partial\overline{z}_{j}}\mathrm{d}z_{i}\otimes\mathrm{d}\overline{z}_{j}$
where $(z_{1},\cdots,z_{n})$ is a local holomorphic coordinate system. Let $(x_{1},y_{1},\cdots,x_{n},y_{n})$ be the corresponding normal coordinates with $z_{k}=x_{k}+iy_{k}$, $k=1,\cdots,n$. A real-valued function on a Kähler manifold is said to be (strictly) plurisubharmonic, if the Levi form is positive (definite) semi-definite. Since $M$ is Kähler, the Levi form $L_{f}$ of $f$ is expressed in terms of Hessian of $f$ using the real coordinates as
$L_{f}={\mathrm{H}}^{f}\left(\frac{\partial}{\partial x_{1}},\frac{\partial}{\partial x_{1}}\right)+{\mathrm{H}}^{f}\left(\frac{\partial}{\partial y_{1}},\frac{\partial}{\partial y_{1}}\right).$
Therefore if $f$ is a strictly convex $C^{2}$-function, then $f$ is strictly plurisubharmonic. As an application to Corollary 3.2, using [24, Lemma 4.2 (i)], we show the following volume growth comparison for metric balls.
Corollary 3.3.
If $M$ as in Theorem 3.1 is also Kähler of (complex) dimension $m$, then $M$ admits a smooth strictly plurisubharmonic function with bounded gradient. Moreover, for each $p\in M$ there exists a constant $C_{2}(p)>0$ such that ,
$V_{p}(r)\geq C_{2}(p)r^{m},\quad r\geq 1$
where $V_{p}(r)$ is the volume of a metric ball of radius $r$ about $p$ in $M$.
4. Spectral theoretic applications
We begin with a sketch of the basic functional analytic setup concerning the spectrum of the Laplacian following [4]. Let $M$ be a complete Riemannian manifold. Then the (positive) Laplacian $\Delta$ is an essentially self-adjoint operator with the space of compactly supported smooth functions on $M$ as the domain. We also denote its unique self-adjoint extension, having the domain $\mathcal{D}(\Delta):=\{f\in L^{2}(M):\Delta f\in L^{2}(M)\}$, by $\Delta$. Using the functional calculus for $\Delta$ we have the family of spectral projection operators $\prod_{E}:=\chi_{E}(\Delta)$ for each Borel subset $E$ of ${\mathbb{R}}$ referred to as the resolution of $\Delta$. The support of these projections and the spectrum of $\Delta$ are related in the following way. We say that $\lambda\in{\mathbb{C}}$ is in the spectrum $\sigma(\Delta)$ of $\Delta$ if and only if $\prod_{(\lambda-\epsilon,\lambda+\epsilon)}\neq 0$ for all $\epsilon>0$. If $\prod_{\{\lambda\}}\neq 0$, then $\lambda$ is an eigenvalue of $\Delta$ and the range of $\prod_{\{\lambda\}}$ is the corresponding eigenspace. The spectral theorem gives a unitary equivalence between $\Delta$ and a multiplication operator on a suitable Hilbert space of the form $\mathrm{L}^{2}(X,\mu)$, where $\mu$ is a $\sigma-$ finite Borel measure on $X$. By the Lebesgue decomposition, $\mu$ decomposes as $\mu=\mu_{pp}+\mu_{ac}+\mu_{sc}$ into pure point measure, absolutely continuous and singularly continuous measures with respect to the Lebesgue measure, and consequently we have that
(4.1) $\mathrm{L}^{2}(X,\mu)=\mathrm{L}^{2}(X,\mu_{pp})\oplus\mathrm{L}^{2}(X,\mu_{ac})\oplus\mathrm{L}^{2}(X,\mu_{sc}).$
Here we take the closure for the set of all eigenvalues as it is not necessarily closed. We refer to $\sigma_{ac}(\Delta)$ and $\sigma_{sc}(\Delta)$ as the absolutely continuous and singularly continuous parts of $\sigma(\Delta)$ respectively and their union as the continuous part $\sigma_{cont}(\Delta)$ of $\sigma(\Delta)$.
In order to analyse the nature of the spectrum of the Laplacian, we invoke [28, Theorem 2]. This result demands boundedness of the gradient and of the Laplacian of a strictly convex function as well as the additional constraint of nonpositivity on the bilaplacian and in turn yields the absolute continuity of the spectrum. Consequently, it enables us to exclude the existence of the point spectrum and the singularly continuous spectrum. As pointed out in [28], the seemingly unnatural additional condition that $\Delta^{2}F\leq 0$ is, in fact, necessary. Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed there. The proof of absolute continuity of the spectrum is then derived from an estimate implementing Kato’s theory of $H-$smooth operators in Hilbert spaces. We shall observe below that the condition on the bilaplacian is satisfied for $M$ as in Theorem 3.1 which are in addition asymptotically harmonic.
@@ 4. Spectral theoretic applications
We first briefly discuss some facts regarding the bottom $\lambda_{0}(M)$ of the spectrum for asymptotically harmonic manifolds. For a complete noncompact Riemannian manifold $M$ it is characterized by [6, §4.4]
(4.2) $\lambda_{0}(M)=\inf\limits_{f\neq 0}\frac{\int\limits_{M}|\nabla f|^{2}}{\int\limits_{M}|f|^{2}}$
@@
where $f$ varies over the completion $\mathfrak{H}(M)$ of $C_{c}^{\infty}(M)$ with respect to the norm $\left\lVert f\right\rVert=\left\lVert f\right\rVert_{\rm{L}^{2}(M)}+\left\lVert\nabla f\right\rVert_{\rm{L}^{2}(M)}$. Let $|\Omega|$ denote the volume of a submanifold $\Omega$ in $M$. Then the Cheeger’s constant of $M$ is given by [6, §4.3]
@@
where the infimum varies over all open, connected submanifolds $\Omega$ of $M$ with compact closure and smooth boundary $\partial\Omega$. The importance of this global geometric invariant is illustrated by the following fundamental inequality relating it to the global analytic invariant $\lambda_{0}(M)$ [6, §4.3, Theorem 3],
(4.3) $\frac{h_{Cheeger}^{2}(M)}{4}\leq\lambda_{0}(M).$
@@
Let $M$ be an asymptotically harmonic manifold with horospheres of mean curvature $h$ i.e. $\Delta b_{v}=h$ for all $v\in SM$. Then we have $h_{Cheeger}=h$ and $\lambda_{0}(M)=h^{2}/4$.
@@
(4.4) $h\leq\inf\limits_{\Omega\subseteq M}\frac{|\partial\Omega|}{|\Omega|}=h_{Cheeger}$
where the infimum varies over $\Omega$ as above in $M$. Let $o\in M$ be arbitrary. Since $M$ is simply connected without conjugate points the exponential at $o$ is a diffeomorphism. Let $d(o,\cdot):M\to{\mathbb{R}}$ denote the distance from $o$. By taking the sequence of functions $f_{n}(x)=e^{-\frac{1}{2}(h+\frac{1}{n})d(o,x)}$ in $\mathfrak{H}(M)$, $n\in{\mathbb{N}}$, we see from (4.2) that
(4.5) $\lambda_{0}(M)\leq\frac{h^{2}}{4}$
Combining (4.3), (4.4), (4.5) we have
(4.6) $h=h_{Cheeger}\quad\text{and}\quad\lambda_{0}(M)=\frac{h^{2}}{4}.$
We remark here that for the proposition above to hold it suffices that there is a $v\in SM$ such that the corresponding horospheres have constant mean curvature.
Lemma 4.2.
Proof.
$\sum\limits_{i=1}^{n}{\mathrm{H}}^{F}(p)(E_{i}(p),E_{i}(p))$ $=\sum\limits_{i=1}^{n}\int\limits_{M(\infty)}{\mathrm{H}}^{b_{v}}(p)(E_{i}(p),E_{i}(p))~{}\mathrm{d}\eta(v)$
$=\int\limits_{M(\infty)}\sum\limits_{i=1}^{n}{\mathrm{H}}^{b_{v}}(p)(E_{i}(p),E_{i}(p))~{}\mathrm{d}\eta(v)$
$\Delta F(p)$ $=\int\limits_{M(\infty)}\Delta b_{v}(p)~{}\mathrm{d}\eta(v)=h$
and the lemma follows. ∎
Theorem 4.3.
If $M$ as in Theorem 3.1 is also asymptotically harmonic with $h$ as the constant mean curvature of the horospheres, then the spectrum of the Laplacian is purely absolutely continuous and is contained in $[\frac{h^{2}}{4},\infty)$.
Proof.
The absolute continuity of the spectrum follows by combining [28, Theorem 2] with Lemma 4.2. That the spectrum of Laplacian is contained in $[\frac{h^{2}}{4},\infty)$ follows from (4.6). ∎
Next we briefly describe harmonic manifolds, only mentioning aspects which are needed in the sequel and to set notation, and we refer to [1, Chapter $6$], and to the comprehensive exposition [21] for the details. A function $f$ on a Riemannian manifold $(M,g)$ is said to be radial around a point $o$ if $f$ is constant on geodesic spheres $S_{o}(r)$ of radius $r$ centered at $o$. $M$ is said to be harmonic if the volume density in normal coordinates centered at $o$ is a radial function. Given a continuous function $f$, its radialisation about $o$ given by
@@
Thus simply connected noncompact nonflat harmonic manifolds necessarily have constant negative Ricci curvature. For every $v\in SM$, the Busemann function $b_{v}$ satisfies $\Delta b_{v}=h$ (see for instance [5, §2, p. 33]) and so it follows that every harmonic manifold is also asymptotically harmonic. Consequently we obtain the following corollary to Theorem 4.3.
@@
Let $M$ be a complete simply connected noncompact nonflat harmonic manifold $M$ without focal points. Then the spectrum of the Laplacian is purely absolutely continuous and is contained in $[h^{2}/4,\infty)$.
We now proceed to show that the spectrum is actually equal to the interval $[h^{2}/4,\infty)$, by making use of the unitarity of the spherical Fourier transform for radial functions on harmonic manifolds of purely exponential volume growth studied in [3]. Notice that in [3] unitarity is shown only in dimensions greater than $5$. However, the only simply connected noncompact nonflat harmonic manifolds of dimension less than equal to $5$ are Riemannian symmetric spaces of noncompact type. This case is handled in Theorem 4.8 below. First, we observe from [21, Theorems 32, 45] that, noncompact harmonic manifolds without focal points fall into the class of noncompact harmonic manifolds of purely exponential volume growth. Let $o$ be an arbitrary but fixed point in $M$. Then for $f\in C_{c}^{\infty}(M)$ which is radial about $o$, the spherical Fourier transform ${\mathcal{F}}^{o}f$ is given by

### v2

@@ 4. A construction using Busemann functions and its applications
Theorem 4.1.
Let $M$ be simply connected Riemannian manifold with no focal points and negative Ricci curvature. Then $F(p)=\displaystyle{\int\limits_{M(\infty)}b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)}$ is a strictly convex function whose gradient is bounded.
Proof.
Let $o$ be an arbitrary but fixed point in $M$. Let $\{v_{k}\}$ be a countable dense subset in $S_{o}M$. Consider the Dirac measures $\delta_{v_{k}}$ associated to each $v_{k}$. Then ${\eta=\sum\limits_{k=1}^{\infty}\frac{1}{2^{k}}\delta_{v_{k}}}$ defines a probability measure on $S_{o}M$. We identify $S_{o}M$, $M(\infty)$ and $\partial^{o}M$ via the respective homeomorphisms. By dominated convergence theorem, the gradient of $F$ at $p$ is given by $\nabla F(p)=\displaystyle{\int\limits_{M(\infty)}\nabla b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)}$. Clearly,
@@
$\left\lVert\nabla F(p)\right\rVert=\left\lVert\int\limits_{M(\infty)}\nabla b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)\right\rVert\leq\int\limits_{M(\infty)}\left\lVert\nabla b_{v}(p)\right\rVert\penalty 10000\ \mathrm{d}\eta(v)=1,\quad\text{as}\penalty 10000\ \left\lVert\nabla b_{v}\right\rVert=1$
i.e., $F$ has bounded gradient. In particular, $F$ is Lipschitz with Lipschitz constant $1$. For $p\in M$ and $w\in S_{p}M$, using (3) and that $\mathrm{Hess}\penalty 10000\ b_{v}$ is positive semidefinite we have
$\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v)$ $=\left|\int\limits_{M(\infty)}{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}\penalty 10000\ \mathrm{d}\eta(v)\right|$
$\leq\int\limits_{M(\infty)}|{\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w},{w\rangle}|\penalty 10000\ \mathrm{d}\eta(v)$
$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w\right\rVert\left\lVert w\right\rVert\penalty 10000\ \mathrm{d}\eta(v)\quad\text{(using Cauchy-Schwarz)}$
$\leq\int\limits_{M(\infty)}\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\penalty 10000\ \mathrm{d}\eta(v)\quad(D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\penalty 10000\ \text{is a bounded linear map for}\penalty 10000\ v\in S_{p}M).$
@@
The restriction of the continuous map (§2) $w\mapsto\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert$ on $SM$ to $S_{p}M$ shows that there exists some $L>0$ such that $\left\lVert D_{w}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $w\in S_{p}M$. In particular, $\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)\right\rVert\leq L$ for all $v\in S_{o}M$. So another application of the dominated convergence theorem shows that
$\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)=\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v).$
@@
The convexity of $F$ follows from the convexity of the Busemann functions. We now claim that $F$ is, in fact, strictly convex. Suppose there exists $p\in M$ and $0\neq w\in T_{p}M$ such that $\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)=0$ then
$\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)\penalty 10000\ \mathrm{d}\eta(v)=0$
@@
Since $v\mapsto\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)$ is a nonnegative continuous function in $v$, we have $\mathrm{Hess}\penalty 10000\ b_{v}(p)(w,w)=0$ for all $v\in S_{o}M$. Using (3),
@@
(6) ${\langle D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}},{D_{\nabla b_{v}(p)}(0)w^{\perp}\rangle}=0,\quad\text{for all}\penalty 10000\ v\in S_{o}M.$
@@
It is known from [Esc77, §5] that for $v\in SM$ and all parallel vector fields $x$ normal to $\gamma_{v}$, $\frac{d}{dt}(\left\lVert D_{v}(t)x\right\rVert^{2})\leq 0$ i.e., the function $f_{v,x}(t)=\frac{1}{2}\left\lVert D_{v}(t)x\right\rVert^{2}$ is $C^{2}$ and monotonically decreasing. From (6) we see that for each $v\in S_{o}M$, $f_{\nabla b_{v}(p),w^{\perp}}$ has a critical point at 0. We claim that $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. If not, then 0 is either a local minimum or local maximum. Suppose 0 is a local minimum then there exists a $\delta$-neighbourhood of 0 such that $f_{\nabla b_{v}(p),w^{\perp}}(0)\leq f_{\nabla b_{v}(p),w^{\perp}}(t)$ for all $t\in(-\delta,\delta)$. Since $f$ is monotonically decreasing we have $f_{\nabla b_{v}(p),w^{\perp}}(t)=f_{\nabla b_{v}(p),w^{\perp}}(0)$ for all $t\in(0,\delta)$. Hence $f_{\nabla b_{v}(p),w^{\perp}}$ is constant on $(0,\delta)$. Observe that $f_{\nabla b_{v}(p),w^{\perp}}$ is $C^{2}$ and so $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=\lim\limits_{t\to 0_{+}}f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(t)=0$, a contradiction. A similar argument could be carried out to show that 0 cannot be a local maximum of $f_{\nabla b_{v}(p),w^{\perp}}$. Hence 0 is a degenerate critical point of $f_{\nabla b_{v}(p),w^{\perp}}$ i.e, $f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)=0$. Then the Jacobi equation for $D_{\nabla b_{v}(p)}$,
@@
$0$ $=f_{\nabla b_{v}(p),w^{\perp}}^{{}^{\prime\prime}}(0)$
@@
(7) $\implies K(\nabla b_{v}(p),w^{\perp})$ $=\left\lVert D_{\nabla b_{v}(p)}^{{}^{\prime}}(0)w^{\perp}\right\rVert^{2}\geq 0,\quad\text{for all $v\in S_{o}M$}$
@@
Now for $\{x_{1},\cdots,x_{n-1}\}$ orthonormal vectors in $T_{p}M$ such that $x_{i}\perp w$, $Ricc(w,w)=\sum\limits_{i=1}^{n}K(x_{i},w)\geq 0$. This is a contradiction to the hypothesis that $M$ has negative Ricci curvature. Hence $\mathrm{Hess}\penalty 10000\ ^{F}(p)(w,w)>0$ for every $p\in M$ and every $0\neq w\in T_{p}M$ proving the strict convexity of $F$. ∎
As indicated at the beginning of the section, we will be concerned with the spectrum on asymptotically harmonic manifolds. Recall that these are simply connected Riemannian manifolds without conjugate points such that horospheres are of constant mean curvature $h\geq 0$. In particular, we answer more refined questions of whether there is a nontrivial point spectrum or singularly continuous spectrum on these manifolds. The main tool here is an implementation of Kato’s theory of $H$-smooth operators on complete Riemannian manifolds due to [Xav88, Theorem 2]. When it can be carried out, it enables one to exclude the existence of the point spectrum and the singularly continuous spectrum. As noted in [Xav88], the seemingly unnatural additional condition $\Delta^{2}F\leq 0$ that is demanded is, in fact, necessary. Indeed, if this condition on the bilaplacian fails, Hadamard manifolds with eigenvalues in the spectrum are constructed there. We shall observe below that the condition on the bilaplacian is satisfied for $M$ as in Theorem 4.1 which are in addition asymptotically harmonic.
Lemma 4.2.
Proof.
$\sum\limits_{i=1}^{n}\mathrm{Hess}\penalty 10000\ ^{F}(p)(E_{i}(p),E_{i}(p))$ $=\sum\limits_{i=1}^{n}\int\limits_{M(\infty)}\mathrm{Hess}\penalty 10000\ b_{v}(p)(E_{i}(p),E_{i}(p))\penalty 10000\ \mathrm{d}\eta(v)$
$=\int\limits_{M(\infty)}\sum\limits_{i=1}^{n}\mathrm{Hess}\penalty 10000\ b_{v}(p)(E_{i}(p),E_{i}(p))\penalty 10000\ \mathrm{d}\eta(v)$
$\Delta F(p)$ $=\int\limits_{M(\infty)}\Delta b_{v}(p)\penalty 10000\ \mathrm{d}\eta(v)=h$
and the lemma follows. ∎
As an immediate consequence we have the following theorem applying [Xav88, Theorem 2].
Theorem 4.3.
If $M$ as in Theorem 4.1 is also asymptotically harmonic with $h$ as the constant mean curvature of the horospheres, then the spectrum of the Laplacian is purely absolutely continuous. Moreover, the spectrum is contained in $\frac{h^{2}}{4},\infty)$.
Proof.
The absolute continuity of the spectrum follows by combining [Xav88, Theorem 2] with Lemma 4.2. That the spectrum of Laplacian is contained in $\frac{h^{2}}{4},\infty)$ follows from the discussion below. ∎
@@
We first briefly discuss some facts regarding the bottom $\lambda_{0}(M)$ of the spectrum for general asymptotically harmonic manifolds. For a complete noncompact Riemannian manifold $M$ it is characterized by [Cha84, §4.4]
(8) $\lambda_{0}(M)=\inf\limits_{f\neq 0}\frac{\int\limits_{M}|\nabla f|^{2}}{\int\limits_{M}|f|^{2}}$
@@
where $f$ varies over the completion $\mathfrak{H}(M)$ of $C_{c}^{\infty}(M)$ with respect to the norm $\left\lVert f\right\rVert=\left\lVert f\right\rVert_{\rm{L}^{2}(M)}+\left\lVert\nabla f\right\rVert_{\rm{L}^{2}(M)}$. Let $|\Omega|$ denote the volume of a submanifold $\Omega$ in $M$. Then the Cheeger’s constant of $M$ is given by
@@
where the infimum varies over all open, connected submanifolds $\Omega$ of $M$ with compact closure and smooth boundary $\partial\Omega$. The importance of this global geometric invariant is illustrated by the following fundamental inequality relating it to the global analytic invariant $\lambda_{0}(M)$ [Cha84, §4.3, Theorem 3],
(9) $\frac{h_{Cheeger}^{2}(M)}{4}\leq\lambda_{0}(M).$
@@
Let $M$ be an asymptotically harmonic manifold with horospheres of mean curvature $h$ i.e., $\Delta b_{v}=h$ for all $v\in SM$. Then we have $h_{Cheeger}=h$ and $\lambda_{0}(M)=h^{2}/4$.
@@
(10) $h\leq\inf\limits_{\Omega\subseteq M}\frac{|\partial\Omega|}{|\Omega|}=h_{Cheeger}$
where the infimum varies over $\Omega$ as above in $M$. Let $o\in M$ be arbitrary. Since $M$ is simply connected without conjugate points the exponential at $o$ is a diffeomorphism. Let $d(o,\cdot):M\to{\mathbb{R}}$ denote the distance from $o$. By taking the sequence of functions $f_{n}(x)=e^{-\frac{1}{2}(h+\frac{1}{n})d(o,x)}$ in $\mathfrak{H}(M)$, $n\in{\mathbb{N}}$, we see from (8) that
(11) $\lambda_{0}(M)\leq\frac{h^{2}}{4}$
Combining (9), (10), (11) we have
(12) $h=h_{Cheeger}\quad\text{and}\quad\lambda_{0}(M)=\frac{h^{2}}{4}$
Remark 4.5.
We remark here that for the proposition above to hold it suffices that, in the case of simply connected Riemannian manifold without conjugate points, there is a single vector $v\in SM$ such that the corresponding horospheres have constant mean curvature.
We have the following immediate corollary to Theorem 4.3 following [Don81] when the sectional curvature is strictly negative.
Corollary 4.6.
Let $M$ be an asymptotically harmonic Hadamard manifold with strictly negative sectional curvature. Then the spectrum of the Laplacian is purely absolutely continuous and is equal to $\frac{h^{2}}{4},\infty)$.
Observe that for such manifolds the essential spectrum and the absolutely continuous spectrum have been determined in [CS14, Theorem 4.4] and [BMP23, Theorem 1.3] respectively.
Harmonic manifolds are important examples of asymptotically harmonic manifolds. We describe only the aspects of harmonic manifolds which are needed in the sequel and to set notation, and we refer to [Bes78, Chapter $6$], and to the comprehensive exposition [Kni16] for the details. A function $f$ on a Riemannian manifold $(M,g)$ is said to be radial around a point $o$ if $f$ is constant on geodesic spheres $S_{o}(r)$ of radius $r$ centered at $o$. $M$ is said to be harmonic if the volume density in normal coordinates centered at $o$ is a radial function. Given a continuous function $f$, its radialisation about $o$ given by
@@
Thus simply connected noncompact nonflat harmonic manifolds necessarily have constant negative Ricci curvature $\kappa<0$. For every $v\in SM$, the Busemann function $b_{v}$ satisfies $\Delta b_{v}=h$ (see for instance [Bra24, §2, p. 33]) and so it follows that every harmonic manifold is also asymptotically harmonic. Consequently we obtain the following corollary to Theorem 4.3.
@@
Let $M$ be a simply connected noncompact (nonflat) harmonic manifold $M$ without focal points. Then the spectrum of the Laplacian is purely absolutely continuous and is contained in $h^{2}/4,\infty)$.
Remark 4.8.
We reiterate here that for this class of manifolds, the negativity of Ricci curvature is not an additional assumption.
We now proceed to give a novel proof that the spectrum is actually equal to the interval $h^{2}/4,\infty)$. For this, we make use of the recently proved [BKP21] unitarity of the Fourier transform for radial functions on harmonic manifolds of purely exponential volume growth, certain estimates on the $c$-function, and the characterisation of harmonic manifolds that the radialisation operator commutes with the Laplacian to pass from radial $\mathrm{L}^{2}$-functions to all $\mathrm{L}^{2}$-functions. We note that unitarity there is shown only in dimensions greater than $5$. However, the only simply connected noncompact (nonflat) harmonic manifolds of dimension less than equal to $5$ are Riemannian symmetric spaces of noncompact type [BKP21, p.145] and this case is covered by the discussion on symmetric spaces at the end.
First, we observe from [Kni16, Theorems 32, 45] that, noncompact harmonic manifolds without focal points fall into the class of noncompact harmonic manifolds of purely exponential volume growth. Let $o$ be an arbitrary but fixed point in $M$. Then for $f\in C_{c}^{\infty}(M)$ which is radial about $o$, the spherical Fourier transform ${\mathcal{F}}^{o}f$ is given by